# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ VII Mixed-state conversion
VII.1 Distillation of asymmetry
Here, we investigate the scenario where we convert a general state $\rho$ to a fixed pure state $\phi$ that serves as a reference. We define the distillable asymmetry of a general state $\rho$ as
$\mathcal{A}_{\mathrm{d}}(\rho)\coloneqq\sup\{r\geq 0\mid\{\rho^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}\},$ (67)
i.e., $\mathcal{A}_{\mathrm{d}}(\rho)=R(\rho\to\phi)$. As a central quantifier of our analysis, we introduce a $\dim G\times\dim G$ Hermitian matrix $\mathcal{S}^{\rho}$ whose elements are given by
$\left(\mathcal{S}^{\rho}\right)_{\mu\nu}\coloneqq\mathrm{Tr}\left(\rho X_{\mu}(I-\Pi_{\rho})X_{\nu}\right),$ (68)
where $\Pi_{\rho}$ denotes the projector to the support of $\rho$. This quantity is an extension of QGT to a mixed state in that $\mathcal{S}^{\psi}=\mathcal{Q}^{\psi}$ for any pure state $\psi$. As we shall show below, the monotonicity of $\mathcal{S}^{\rho}$ under $G$-covariant channels yields an upper bound of $\mathcal{A}_{\mathrm{d}}(\rho)$, from which we can derive a sufficient condition for the rate to vanish.
We start the analysis by relating $\mathcal{S}^{\rho}$ with the monotone metric with an operator monotone function $f_{q}(x)$ given in Eq. (24). By using the eigenvalue decomposition $|\rho=\sum_{k}p_{k}\ket{k}\bra{k}$ of the state $\rho$, we have
@@ VII.1 Distillation of asymmetry
In particular, when $O=\bm{\gamma}^{\dagger}\bm{X}$ with $\bm{\gamma}\in\mathbb{C}^{\dim G}$, we get
$\lim_{q\to 1^{-}}f_{q}(0)\|\mathrm{i}[\bm{\gamma}^{\dagger}\bm{X},\rho]\|_{f_{q},\rho}^{2}=\bm{\gamma}^{\dagger}\mathcal{S}^{\rho}\bm{\gamma}.$ (71)
We now prove the monotonicity of $\mathcal{S}^{\rho}$ under $G$-covariant operation. Suppose a state $\rho$ is convertible to $\sigma$ via a $G$-covariant channel without error. Multiplying Eq. (32) by $f_{q}(0)=1-q>0$ yields
$f_{q}(0)\|\mathrm{i}[\bm{\gamma}^{\dagger}\bm{X},\rho]\|_{f_{q},\rho}^{2}\geq f_{q}(0)\|\mathrm{i}[\bm{\gamma}^{\dagger}\bm{X}^{\prime},\sigma]\|_{f_{q},\sigma}^{2}.$ (72)
Taking the limit of $q\to 1^{-}$, Eq. (71) implies $\bm{\gamma}^{\dagger}\mathcal{S}^{\rho}\bm{\gamma}\geq\bm{\gamma}^{\dagger}\mathcal{S}^{\sigma}\bm{\gamma}$. Since this inequality holds for any $\bm{\gamma}\in\mathbb{C}^{\dim G}$, we get
$\mathcal{S}^{\rho}\geq\mathcal{S}^{\sigma},$ (73)
meaning that $\mathcal{S}^{\rho}$ is a valid asymmetry monotone for any state $\rho$.
The asymptotic discontinuity of QGT is inherited by $\mathcal{S}^{\rho}$, as they coincide for pure states, implying that extending its monotonicity to an asymptotic setting requires careful consideration. Nevertheless, as an extension of Proposition 5, an upper bound of the distillable asymmetry can be proven by using Lemma 4.
@@
Let $U,U^{\prime}$ be projective unitary representations of a Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. For any state $\rho$ and pure state $\phi$, it holds
$\sup\{r\geq 0\mid\forall g\in G,\,\mathcal{S}^{\mathcal{U}_{g}(\rho)}\geq r\mathcal{Q}^{\mathcal{U}^{\prime}_{g}(\phi)}\}\geq\mathcal{A}_{\mathrm{d}}(\rho).$ (74)
@@
Suppose $\{\rho^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$ for $r\geq 0$. From Eq. (36) and Lemma 4, as an extension of Eq. (38), we get
$f_{q}(0)\|\mathrm{i}[\bm{\gamma}^{\dagger}\bm{X},\rho]\|_{f_{q},\rho}^{2}\geq\frac{M}{N}\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}-\frac{M}{N}h(\epsilon)+\frac{1}{N}o(M),$ (75)
@@
$f_{q}(0)\|\mathrm{i}[\bm{\gamma}^{\dagger}\bm{X},\rho]\|_{f_{q},\rho}^{2}\geq r\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}-rh(\epsilon).$ (76)
Since this inequality holds for all sufficiently small $\epsilon>0$, we get $f_{q}(0)\|\mathrm{i}[\bm{\gamma}^{\dagger}\bm{X},\rho]\|_{f_{q},\rho}\geq r\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}$. In the limit of $q\to 1^{-}$, we obtain $\mathcal{S}^{\rho}\geq r\mathcal{Q}^{\phi}$. Repeating this argument for $\mathcal{U}_{g}(\rho)$ and $\mathcal{U}^{\prime}_{g}(\phi)$ instead of $\rho$ and $\phi$, we get
$\forall g\in G,\quad\mathcal{S}^{\mathcal{U}_{g}(\rho)}\geq r\mathcal{Q}^{\mathcal{U}^{\prime}_{g}(\phi)},$ (77)
@@
This result first establishes an upper bound on the distillable asymmetry, which applies to any state $\rho$ and any compact Lie group $G$, and becomes tight when $\rho$ is pure.
We remark that in a special case where $\Pi_{\rho},\bm{\gamma}^{\dagger}\bm{X}]=0$ holds for some $\bm{\gamma}\in\mathbb{C}^{\dim G}$, we have $\bm{\gamma}^{\dagger}\mathcal{S}^{\rho}\bm{\gamma}=0$. As an immediate corollary of the above theorem, we therefore get the following:
@@
If there exists a vector $\bm{\gamma}\in\mathbb{C}^{\dim G}$ satisfying $\Pi_{\rho},\bm{\gamma}^{\dagger}\bm{X}]=0$ and $\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}\neq 0$, then $\mathcal{A}_{\mathrm{d}}(\rho)=0$.
This corollary implies that for a typical state with full rank, $\mathcal{A}_{\mathrm{d}}(\rho)=0$ holds for any asymmetric pure state $\phi$. This result is consistent with a result in prior research [23] on time-translation asymmetry, where the distillable coherence, defined as the optimal conversion rate from a state $\rho$ to a pure state $\phi$ given by a superposition of the ground and excited state of a Hamiltonian $H$, is shown to be zero when $\Pi_{\rho},H]=0$. We remark that although a state $\rho$ satisfying the condition in Corollary 13 has zero distillable asymmetry, it does not necessarily implies that $\rho$ is useless in asymptotic conversion. For example, while a full-rank state has zero distillable asymmetry, a sublinear number of such states may be used to enhance the conversion rate by removing the restriction imposed by symmetry subgroups in asymptotic conversion among pure states, as stated in Theorem 11.
VII.2 Dilution of asymmetry
@@
$\mathcal{A}_{\mathrm{c}}(\rho)\coloneqq\inf\{r\geq 0\mid\{\phi^{\otimes\lceil rN\rceil}\}_{N}\xrightarrow{G\text{-cov.}}\{\rho^{\otimes N}\}_{N}\},$ (78)
i.e., $\mathcal{A}_{\mathrm{c}}(\rho)=1/R(\rho\to\phi)$, where $1/\infty$ and $1/0$ are formally regarded as 0 and $\infty$, respectively. We remark that Eq. (9) implies $\mathcal{A}_{\mathrm{c}}(\rho)<\infty$ only if $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$. Thus, we only investigate $\mathcal{A}_{\mathrm{c}}(\rho)$ for $\rho$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$.
@@
$r_{i}=2^{\min_{j}D_{\max}\left(\mathcal{Q}^{\mathcal{U}_{g_{j}}^{\prime}(\psi_{i})}\middle\|\mathcal{Q}^{\mathcal{U}_{g_{j}}(\phi)}\right)},$ (82)
where $g_{j}\in G_{j}$ is a representative element of each connected component $G_{j}$ of $G$.
@@
In contrast to the $U(1)$ case, a decomposition into $\rho=\sum_{i=1}^{k-1}p_{i}\psi_{i}+p_{\mathrm{s}}\rho_{\mathrm{s}}$ with symmetric $\rho_{\mathrm{s}}$ and pure states $\{\psi_{i}\}_{i}$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\psi_{i})$ does not always exist for a general Lie group $G$. See Appendix E for a concrete example for $G=U(4)$. In such cases, Eq. (80) does not provide a meaningful upper bound on the asymmetry cost. However, by carefully refining the estimate-and-convert approach introduced in Sec. V.3, we find that this limitation due to the symmetry subgroups can be circumvented. Specifically, in Appendix F, we show the following:
@@
For a reference state $\phi$, consider a state $\rho$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$. The asymmetry cost of a state $\rho$ is upper bounded as follows:

### v4

@@ V Mixed-state conversion
V.1 Distillation of asymmetry
Here, we investigate the scenario where we convert a general state $\rho$ to a fixed pure state $\phi$ that serves as a reference. We define the distillable asymmetry of a general state $\rho$ as $\mathcal{A}_{\mathrm{d}}(\rho)\coloneqq R(\rho\to\phi)$, i.e.,
$\mathcal{A}_{\mathrm{d}}(\rho)=\sup\{r\geq 0\mid\{\rho^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}\}.$ (42)
As a central quantity in our analysis, we introduce a $\dim G\times\dim G$ Hermitian matrix $\mathcal{Q}^{\rho}$ whose elements are given by
$\left(\mathcal{Q}^{\rho}\right)_{\mu\nu}\coloneqq\mathrm{Tr}\left(\rho X_{\mu}(I-\Pi_{\rho})X_{\nu}\right),$ (43)
where $\Pi_{\rho}$ denotes the projector to the support of $\rho$. This quantity is an extension of QGT, originally defined for pure states in Eq. (7), to mixed states. As we shall show below, the monotonicity of $\mathcal{Q}^{\rho}$ under $G$-covariant channels yields an upper bound of $\mathcal{A}_{\mathrm{d}}(\rho)$, from which we can derive a sufficient condition for the rate to vanish.
We start the analysis by relating $\mathcal{Q}^{\rho}$ with the monotone metric with an operator monotone function $f_{q}(x)$ given in Eq. (25). By using the eigenvalue decomposition $|\rho=\sum_{k}p_{k}\ket{k}\bra{k}$ of the state $\rho$, we have
@@ V.1 Distillation of asymmetry
In particular, when $O=\gamma^{\dagger}X$ with $\gamma\in\mathbb{C}^{\dim G}$, we get
$\lim_{q\to 1^{-}}f_{q}(0)\|\mathrm{i}[\gamma^{\dagger}X,\rho]\|_{f_{q},\rho}^{2}=\gamma^{\dagger}\mathcal{Q}^{\rho}\gamma.$ (46)
We now prove the monotonicity of $\mathcal{Q}^{\rho}$ under $G$-covariant operation. Suppose a state $\rho$ is convertible to $\sigma$ via a $G$-covariant channel without error. Multiplying Eq. (30) by $f_{q}(0)=1-q>0$ yields
$f_{q}(0)\|\mathrm{i}[\gamma^{\dagger}X,\rho]\|_{f_{q},\rho}^{2}\geq f_{q}(0)\|\mathrm{i}[\gamma^{\dagger}X^{\prime},\sigma]\|_{f_{q},\sigma}^{2}.$ (47)
Taking the limit of $q\to 1^{-}$ in Eq. (47), Eq. (46) implies $\gamma^{\dagger}\mathcal{Q}^{\rho}\gamma\geq\gamma^{\dagger}\mathcal{Q}^{\sigma}\gamma$. Since this inequality holds for any $\gamma\in\mathbb{C}^{\dim G}$, we get
$\mathcal{Q}^{\rho}\geq\mathcal{Q}^{\sigma},$ (48)
meaning that $\mathcal{Q}^{\rho}$ is a valid asymmetry monotone for any state $\rho$.
The asymptotic discontinuity of QGT is inherited by $\mathcal{Q}^{\rho}$, as they coincide for pure states, implying that extending its monotonicity to an asymptotic setting requires careful consideration. Nevertheless, as an extension of Eq. (38), an upper bound of the distillable asymmetry can be proven by using Eq. (34).
@@
Let $U,U^{\prime}$ be projective unitary representations of a Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. For any state $\rho$ on $\mathcal{H}$ and pure state $\phi$ on $\mathcal{H}^{\prime}$, it holds
$\sup\{r\geq 0\mid\forall g\in G,\,\mathcal{Q}^{\mathcal{U}_{g}(\rho)}\geq r\mathcal{Q}^{\mathcal{U}^{\prime}_{g}(\phi)}\}\geq\mathcal{A}_{\mathrm{d}}(\rho).$ (49)
@@
Suppose $\{\rho^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$ for $r\geq 0$. From Eqs. (33) and (34), as an extension of Eq. (35), we get
$f_{q}(0)\|\mathrm{i}[\gamma^{\dagger}X,\rho]\|_{f_{q},\rho}^{2}\geq\frac{M}{N}\gamma^{\dagger}\mathcal{Q}^{\phi}\gamma-\frac{M}{N}h(\epsilon)+\frac{1}{N}o(M),$ (50)
@@
$f_{q}(0)\|\mathrm{i}[\gamma^{\dagger}X,\rho]\|_{f_{q},\rho}^{2}\geq r\gamma^{\dagger}\mathcal{Q}^{\phi}\gamma-rh(\epsilon).$ (51)
Since this inequality holds for all sufficiently small $\epsilon>0$, we get $f_{q}(0)\|\mathrm{i}[\gamma^{\dagger}X,\rho]\|_{f_{q},\rho}\geq r\gamma^{\dagger}\mathcal{Q}^{\phi}\gamma$. In the limit of $q\to 1^{-}$, we obtain $\mathcal{Q}^{\rho}\geq r\mathcal{Q}^{\phi}$. Repeating this argument for $\mathcal{U}_{g}(\rho)$ and $\mathcal{U}^{\prime}_{g}(\phi)$ instead of $\rho$ and $\phi$, we get
$\forall g\in G,\quad\mathcal{Q}^{\mathcal{U}_{g}(\rho)}\geq r\mathcal{Q}^{\mathcal{U}^{\prime}_{g}(\phi)},$ (52)
@@
This result first establishes an upper bound on the distillable asymmetry by using a geometric quantity, $\mathcal{Q}$, i.e., the QGT and its extension to mixed states, which applies to any state $\rho$ and any Lie group $G$. From Theorem 1, the inequality (49) becomes tight when $\rho$ is pure, and $G$ is compact.
We remark that in a special case where $\Pi_{\rho},\gamma^{\dagger}X]=0$ holds for some $\gamma\in\mathbb{C}^{\dim G}$, we have $\gamma^{\dagger}\mathcal{Q}^{\rho}\gamma=0$. As an immediate corollary of the above theorem, we therefore get the following:
@@
If there exists a vector $\gamma\in\mathbb{C}^{\dim G}$ satisfying $\Pi_{\rho},\gamma^{\dagger}X]=0$ and $\gamma^{\dagger}\mathcal{Q}^{\phi}\gamma\neq 0$, then $\mathcal{A}_{\mathrm{d}}(\rho)=0$.
This corollary implies that for a typical state with full rank, where $\Pi_{\rho}=I$, $\mathcal{A}_{\mathrm{d}}(\rho)=0$ holds for any asymmetric pure state $\phi$. This result is consistent with a result in prior research [23] on time-translation asymmetry, where the distillable coherence, defined as the optimal conversion rate from a state $\rho$ to a pure state $\phi$ given by a superposition of the ground and excited state of a Hamiltonian $H$, is shown to be zero when $\Pi_{\rho},H]=0$. We remark that although a state $\rho$ satisfying the condition in Corollary 6 has zero distillable asymmetry, it does not necessarily implies that $\rho$ is useless in asymptotic conversion. For example, while a full-rank state has zero distillable asymmetry, a sublinear number of such states may be used to enhance the conversion rate by removing the restriction imposed by symmetry subgroups in asymptotic conversion among pure states, as shown in Appendix K.
V.2 Dilution of asymmetry
@@
$\mathcal{A}_{\mathrm{c}}(\rho)\coloneqq\frac{1}{R(\phi\to\rho)},$ (53)
where $1/\infty$ and $1/0$ are formally regarded as 0 and $\infty$, respectively. From the definition of conversion rate $R(\phi\to\rho)$, we also obtain an equivalent expression
$\mathcal{A}_{\mathrm{c}}(\rho)=\inf\{r\geq 0\mid\{\phi^{\otimes\lceil rN\rceil}\}_{N}\xrightarrow{G\text{-cov.}}\{\rho^{\otimes N}\}_{N}\}.$ (54)
We remark that Eq. (10) implies $\mathcal{A}_{\mathrm{c}}(\rho)<\infty$ only if $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$. Thus, we only investigate $\mathcal{A}_{\mathrm{c}}(\rho)$ for $\rho$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$.
@@
$r_{i}=2^{D_{\max}\left(\mathcal{Q}^{\psi_{i}}\middle\|\mathcal{Q}^{\phi}\right)}.$ (58)
@@
In contrast to the $U(1)$ case, a decomposition into $\rho=\sum_{i=1}^{k-1}p_{i}\psi_{i}+p_{\mathrm{s}}\rho_{\mathrm{s}}$ with symmetric $\rho_{\mathrm{s}}$ and pure states $\{\psi_{i}\}_{i}$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\psi_{i})$ does not always exist for a general Lie group $G$. See Appendix E for a concrete example for $G=U(4)$. In such cases, Eq. (56) does not provide a meaningful upper bound on the asymmetry cost. However, by carefully refining the estimate-and-convert approach introduced in Sec. IV.2, we find that this limitation due to the symmetry subgroups can be circumvented. Specifically, in Appendix F, we show the following:
@@
For a reference pure state $\phi$, consider a state $\rho$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$. The asymmetry cost of a state $\rho$ is upper bounded as follows: